OpenStudy (anonymous):

I have to show that y=x and so far I've simplified it to this: __(in comments)__. Please help, I don't know where to go from here. Thanks :)

OpenStudy (anonymous):

\[x ^{3}+xy+y ^{2}=-1\]

ganeshie8 (ganeshie8):

what exactly are you trying to do here ?

OpenStudy (anonymous):

show that x=y

jimthompson5910 (jim_thompson5910):

you wrote "so far I've simplified it to this" so I'm assuming there's an initial problem that you didn't post? or is \(x ^{3}+xy+y ^{2}=-1\) the original problem?

OpenStudy (anonymous):

oh...sorry, we have to show (from the definition) that \[f(x) = x ^{3} + x \] is a 1-1 function.. so I let \[x ^{3}+x=y^2+y\] then said... \[x ^{3}-y ^{3}=y-x\] Then... \[(x-y)(x ^{2}+xy+y ^{2}) = y-x\] then... \[x ^{3}+xy+y ^{2}=\frac{ -(x-y) }{ (x-y) }\]

OpenStudy (anonymous):

one of your powers is off i think

OpenStudy (anonymous):

argh! sorry...I just realised that... it should be x^2+xy+y^2=-1

ganeshie8 (ganeshie8):

\[\large (x-y)(x ^{2}+xy+y ^{2}) = y-x \iff x^{\color{Red}{2}}+xy+y^2 = -1\]

ganeshie8 (ganeshie8):

when \(x\ne y\)

ganeshie8 (ganeshie8):

otherwise you won't be able to divide x-y and cancel it

OpenStudy (anonymous):

oh..well if I can't cancel then how do I go about it.

ganeshie8 (ganeshie8):

Notice that x^2+xy+y^2 = (x+y)^2 - xy familiar with AM-GM inequality ?

OpenStudy (anonymous):

no...sorry

ganeshie8 (ganeshie8):

Okay, what we're trying basically is : \[\large (x-y)(x ^{2}+xy+y ^{2}) = y-x \\ \large (x-y)( x^{\color{Red}{2}}+xy+y^2 +1)= 0 \]

ganeshie8 (ganeshie8):

By zero product property, either the first factor or the second factor or both can equal 0. o if you prove that the second factor can never equal 0, you're done.

OpenStudy (anonymous):

wait...how did you get that second line...to equal zero. I don't see what you did.

ganeshie8 (ganeshie8):

factored `x-y`

ganeshie8 (ganeshie8):

\[\large (x-y)(x ^{2}+xy+y ^{2}) = y-x \\ \large (x-y)(x ^{2}+xy+y ^{2})+(x-y) =0 \\ \large (x-y)( x^{\color{Red}{2}}+xy+y^2 +1)= 0\]

OpenStudy (anonymous):

oh thanks! so.. (x-y) and (x^2+xy+y^2 +1) both have to equal zero for x=y right?

ganeshie8 (ganeshie8):

Not exactly

ganeshie8 (ganeshie8):

tell me this : 2*3 = 0 is that be true ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

what about this : a*3 = 0 can this be true ever ?

OpenStudy (anonymous):

oh so is either (x-y) or ('the other factor') equal to 0.

ganeshie8 (ganeshie8):

you might be seeing it already a*b = 0 means, either a or b has to be 0

ganeshie8 (ganeshie8):

exactly!

OpenStudy (anonymous):

awesome! Thanks so much for your help :)

ganeshie8 (ganeshie8):

if we could show that "the other factor" is never 0, then that forces the first factor to equal 0

ganeshie8 (ganeshie8):

can you show that "the other factor" is never 0 ?

ganeshie8 (ganeshie8):

it is easy to show using AM-GM inequality, but since you don't want to use it, you need to find some other means

OpenStudy (anonymous):

umm...well \[(x+y) = + or - \sqrt{-1-xy}\] does that show anything?

ganeshie8 (ganeshie8):

Here is how I would work it \[\large (x-y)(x ^{2}+xy+y ^{2}) = y-x \\ \large (x-y)(x ^{2}+xy+y ^{2})+(x-y) =0 \\ \large (x-y)( x^{\color{Red}{2}}+xy+y^2 +1)= 0 \\\large (x-y)=0 ~\lor ~( x^{\color{Red}{2}}+xy+y^2 +1)= 0\] the minimum value of \(\large x^2 + xy+y^2+1 \) is \(\large 1\), so it can never equal \(\large 0\). That forces the first factor to equal \(0\) : \[\large x-y = 0 \implies x = y\]

jimthompson5910 (jim_thompson5910):

When you say "the minimum value of x^2 + xy+y^2+1 is 1", can't it go lower? x could be some very large negative number while y is positive and I can see it dipping below 1.

ganeshie8 (ganeshie8):

By AM-GM inequality \(\large (x+y)^2 \ge 4xy\)

jimthompson5910 (jim_thompson5910):

oh i guess the x^2 counterbalances that

ganeshie8 (ganeshie8):

\[\large x^2 + xy+y^2 = (x+y)^2 - xy \ge 0\] But the OP is not familiar with this inequality so I thought i better skip the explanation and let him figure out his own way of showing this :)

jimthompson5910 (jim_thompson5910):

what I'd do is solve for x in x^2 + xy + y^2 + 1 = 0 by using the quadratic formula x^2 + xy + y^2 + 1 = 0 or x^2 + y*x + y^2 + 1 = 0 is in the form ax^2 + bx + c = 0 where a = 1 b = y c = y^2 + 1

jimthompson5910 (jim_thompson5910):

actually, you can focus on the discriminant formula D = b^2 - 4ac if D < 0, then there won't be any real-number solutions to the quadratic

ganeshie8 (ganeshie8):

that works, its just a quadratic in two variables after all

OpenStudy (anonymous):

oh ok cool!